Westerville Company reported the following results from last year’s operations: Sales $ 1,200,000 Variable expenses 320,000 Contribution margin 880,000 Fixed expenses 640,000 Net operating income $ 240,000 Average operating assets $ 600,000 At the beginning of this year, the company has a $150,000 investment opportunity with the following cost and revenue characteristics: Sales $ 240,000 Contribution margin ratio 50 % of sales Fixed expenses $ 84,000 The company’s minimum required rate of return is 15%.

5. What is the turnover related to this year’s investment opportunity?

6. What is the ROI related to this year’s investment opportunity?

7. If the company pursues the investment opportunity and otherwise performs the same as last year, what margin will it earn this year? (Round your percentage answer to 1 decimal place (i.e .1234 should be entered as 12.3))

8. If the company pursues the investment opportunity and otherwise performs the same as last year, what turnover will it earn this year?

9. If the company pursues the investment opportunity and otherwise performs the same as last year, what ROI will it earn this year? (Round your percentage answer to 1 decimal place (i.e., 0.1234 should be considered as 12.3%.))
Answer #1

Net operating income investment opportunity = (240000*50%-84000) = 36000

5) Turnover = Sales/Average assets = 1200000/600000 = 2

6) ROI = Net operating income/Average assets = 240000/600000 = 40%

7) Margin = 276000/1440000 = 19.2%

8) Turnover = 1440000/750000 = 1.92

9) ROI = 276000/750000 = 36.8%